OpenStudy (anonymous):

solve the equation x^2+3=0

OpenStudy (anonymous):

x^2=-3, so this equation has no solutions, because x^2>=0

OpenStudy (anonymous):

No real solutions that is.

OpenStudy (anonymous):

thank u...

OpenStudy (anonymous):

In imaginary numbers, \(i=\sqrt{-1}\), and hence the solution will be \[x=\sqrt{-3}=i\sqrt{3}\] As pointed out above there are no real solutions to the equation